OpenStudy (bunnyzrockgirl):

Write an equation in point-slope form for the line with the given slope that contains the given point. slope = 4; (3, 8)

OpenStudy (anonymous):

y-8=4(x-3)

OpenStudy (anonymous):

simply plug in for the equation of x-x1=m(slope)(x-x1)

OpenStudy (bunnyzrockgirl):

Thanks!

OpenStudy (anonymous):

yeppers